OpenStudy (anonymous):

Help? Determine whether each matrix has an inverse. If an inverse exists, find it. [ 4,8_-3,-2]

OpenStudy (anonymous):

@phi ?

OpenStudy (fibonaccichick666):

does the det=0?

OpenStudy (anonymous):

no?

terenzreignz (terenzreignz):

\[\huge \left[\begin{matrix}4 & 8 \\ -3 & -2\end{matrix}\right]\] This matrix?

terenzreignz (terenzreignz):

Well, anyway, two things to keep in mind... if the determinant of a matrix is not zero, then it will have an inverse. Second, and importantly, since this is just a 2x2 matrix, there is a straightforward way of getting its inverse... So, suppose we have an (invertible) 2x2 matrix... \[\huge \left[\begin{matrix} a & b \\ c & d\end{matrix}\right]\] Then the quickest way to get its inverse is using this formula... \[\huge \frac1{ad-bc}\left[\begin{matrix} d & -b \\ -c & a\end{matrix}\right]\]

OpenStudy (anonymous):

So [2,3_-4,-8] @terenzreignz

OpenStudy (fibonaccichick666):

no, maybe try using the identity method instead of matrix of cofactors

OpenStudy (anonymous):

Im not sure how to do that?

OpenStudy (fibonaccichick666):

ok so you want to set up a matrix where on the left you have your matrix and then on the right you have the identity matrix

OpenStudy (fibonaccichick666):

ie. \[\huge \left[\begin{matrix} a & b & 1 & 0 \\ c & d & 0 & 1\end{matrix}\right]\]

OpenStudy (fibonaccichick666):

then using row reductions, get the identity on the left instead

OpenStudy (anonymous):

Ok so this does have an inverse??

OpenStudy (fibonaccichick666):

do you really think I'd have you doing all of this if it didn't? And since you told me the \(det \not= 0\) then there does exist an inverse

OpenStudy (anonymous):

the inverse is [-2,3_-8,4]?